On monday morning I had a certain amount of money. By the evening I spent half of want I had pls $3 more: on tuesday I spent half of what I had left plus $3 more: On Wenesday I spent half of what i had plus $3 more and I ended up with only $2. How much cash did I had on Monday morning?
Answer by BrittanyM(80)

In order to solve this problem, we must work backwards.
If on Wednesday, you had $2,
and one day earlier you had double this plus $3,
you would have had $7 on Tuesday.
(this may be represented by 2x + 3, where x=2)
And one day earlier than this, you had double plus $3,
(2x + 3, where x=7)
Monday afternoon, you would have had $17.
Once more, for Monday morning,
(2x + 3, where x=17)
So, Monday morning, you would have had $37.
You may check this by evaluating these values in the problem:
1/2(37-3)=17
1/2(17-3)=7
1/2(7-3)=2